OpenStudy (anonymous):

The difference of two integers is 6. The sum of the same two integers is 42. What is the lesser of the two integers?

OpenStudy (anonymous):

you can probably guess and check for this one or solve \[x-y=6\] \[x+y=42\] by adding the two equations together

OpenStudy (anonymous):

if it was me i would just guess say 20 and 26 were the two numbers, then the total would be 46 not 42 so i guessed too high

OpenStudy (anonymous):

my choices r 18 19 21 23 24

OpenStudy (anonymous):

then you have even fewer guesses to make i tried with the smaller number as 20 that made the larger number 26 since they are 6 apart, but \(20+26=46\) and you want 42, so 20 is too large try a smaller one

OpenStudy (anonymous):

ok